# Phase space analysis of CCDM cosmologies

### v1

@@ III.1.2 Two-fluid system
    When $w<-1$: If the equation of state $w$ of the second fluid lies in the phantom region, we again obtain five critical point which are shown in Table  1, i.e. $A_{0},~{}A_{1},~{}A_{2},~{}A_{3}$ and $A_{4}$. Since the eigenvalues are both negative at $A_{0}$, this is a stable point that exhibits dark-energy dominated accelerated universe. Being two eigenvalues, positive at $A_{1}$, this point is unstable which leads to the matter dominated decelerating phase of the universe. Clearly, the points $A_{0},~{}A_{2}$ properly describe the phase of the universe. Again, as eigenvalues are opposite in sign evaluated at $A_{3},~{}A_{4}$, these two critical points are saddle by nature representing acceleration with completely matter domination. Clearly, $A_{2}$, dark-energy dominated point is unstable because one can easily check that $x^{\prime}$, $z^{\prime}$ are both positive on $z=1$ and $x=0$, respectively. The right plot of Fig. 2 exhibits the phase space behavior where a class of orbits leave the decelerating phase and end in the super accelerating phase which is expected in the phantom cosmology.
@@ III.2 Model: $\Gamma=\Gamma_{0}H$
Here, we consider the particle creation rate, $\Gamma=\Gamma_{0}H$ where $\Gamma_{0}$ is a positive constant. In the following we analyze its effects on the universe evolution for both the systems, i.e. one fluid system and two-fluid system.
III.2.1 One fluid system
This is a special matter creation rate in which the decelerating parameter, $q=\frac{1}{2}(1-\Gamma_{0})$ depends only on the free parameter $\Gamma_{0}$, and hence, one can see that universe experiences an accelerating phase for $\Gamma_{0}>1$ or a decelerating phase for $\Gamma_{0}<1$. Note that for any other matter creation model, $q$ will depend on the $z$ variable. Now, here only $z$ variable given in (24) leads to the one dimensional autonomous dynamical system in the form:
$z^{\prime}=\frac{3}{2}z(1-z)\left(1-\frac{\Gamma_{0}}{3}\right).$ (30)
Clearly we have two critical points, namely, $B^{\prime}_{0}(1,0)$ and $B^{\prime}_{1}(1,1)$, where first and second components of each critical point are related to DM density parameter and values of $z$ variable, respectively. It is obvious from the above dynamical system that $z$ is increasing and decreasing depending on the parameter $\Gamma_{0}<3$ and $\Gamma_{0}>3$, respectively. So, $B^{\prime}_{0}$ shows unstable nature if $\Gamma_{0}<3$ and stable nature if $\Gamma_{0}>3$, while $B^{\prime}_{1}$ gives stable character if $\Gamma_{0}<3$ and unstable character if $\Gamma_{0}>3$. Again, when $\Gamma_{0}=3$, all points in the phase space (i.e. closed interval $[0,1]$) are critical points showing acceleration. All the information about this model are depicted in Table 2 and the vector plots of the system (30) are exhibited in the Fig. 3. Here, both the critical points simultaneously lie either in the accelerating phase (for $\Gamma_{0}>1$) or decelerating phase $(0<\Gamma_{0}<1)$. For $\Gamma_{0}>3$ we also have physical solutions because $q$ is less than $-1$, and $B^{\prime}_{0}(H\longrightarrow+\infty)$ is stable and $B^{\prime}_{1}(H\longrightarrow 0)$ is unstable. Here we can only conclude that our universe ends in an accelerating phase or a decelerating phase for ($1<\Gamma_{0}$, $\Gamma_{0}\neq 3$) or $(0<\Gamma_{0}<1)$, respectively.
III.2.2 Two fluid system
@@
where $\Gamma_{0}$ is a positive constant. The dynamical system is free from singularity. One can clearly see that the above system has four invariant manifolds $x=0,~{}x=1,~{}z=0$ and $z=1$ which make the phase space domain, namely, $R_{2}=[0,1]^{2}$ positively invariant set. In this model the decelerating parameter takes the form $q=\frac{1}{2}\left[1+3w(1-x)-\Gamma_{0}x\right]$ which does not contain the $z$ variable, and for $\left(1-\Gamma_{0}\right)x+(1+3w)(1-x)<0$, the accelerating expansion of the universe is realized. In the Table 2, we properly present the inherent property of the dynamical system (31) (i.e. critical points and different values of the cosmological parameters at these critical points). Since the evolution of $x$ variable significantly depends on the factor $\left(3w+\Gamma_{0}\right)$ for which we carry out the phase space analysis in three possible ways, such as
  1. 1.
    When $03w+\Gamma_{0}<$ 0: In this case, one can notice from Table 2 that there are four isolated critical points $B_{0},~{}B_{1},~{}B_{2},~{}B_{3}$ and two critical lines $B_{4}$(if $w=-1$), $B_{5}$  (if $\Gamma_{0}=3$). Here, at late times, $x$ goes to zero, and then, at late times, we will have $z^{\prime}=\frac{3}{2}z(1-z)(1+w)$. Thus, the following sub-cases will appear:
    1. (a)
      $w<-1:$
      1. i.
@@
        $\Gamma_{0}\neq 3:$ In this situation, the system has only four isolated critical points, namely, $B_{0},~{}B_{1},~{}B_{2},~{}B_{3}$. At late times, $z$ goes to zero, concluding that $B_{0}$ is a global attractor. Again, examining the eigenvalues, one can conclude that $B_{1}$ is saddle for $\Gamma_{0}>3$ and unstable for $\Gamma_{0}<3$, $B_{2}$ is always saddle in nature, $B_{3}$ is unstable if $\Gamma_{0}>3$ and saddle by behavior if $\Gamma_{0}<3$. Here, $B_{0},~{}B_{2}$ always lie in the accelerated region while $B_{1},~{}B_{3}$ remain in the accelerated phase if $\Gamma_{0}>1$. So we can conclude that $B_{0}$ is the late time DE dominated accelerating global attractor and universe undergoes from decelerating to accelerating phase if we choose the model parameter $\Gamma_{0}<1$. This scenario is depicted in the lower right plot of the Fig. 4.
      2. ii.
        $\Gamma_{0}=3:$ Along with two isolated critical points $B_{0},~{}B_{2}$, we have one critical line $B_{5}$, containing the critical points $B_{1},~{}B_{3}$. As one eigenvalue is positive and other is zero, $B_{5}$ behaves like unstable point. All the critical points and line are representing accelerating solutions. Once again, $B_{0}$ is the only late time DE dominated global attractor which we have seen in the previous paragraph.
    2. (b)
      $w=-1:$ When the equation of state of the second fluid coincides with cosmological constant $(w=-1)$, one can get two critical points and one critical line, namely, $B_{1},~{}B_{3}$ and $B_{4}$. Now, we have $z^{\prime}=\frac{3}{2}z(1-z)x\left(1-\frac{\Gamma_{0}}{3}\right)$. Taking into account that $01-\frac{\Gamma_{0}}{3}>$ 0, we deduce that $z$ is an increasing function, concluding that $B_{4}$ behaves like stable in nature. The line $B_{4}$ is always accelerated while $B_{1},~{}B_{3}$ are accelerated for $\Gamma_{0}>1$ otherwise decelerated. Thus, one may conclude that the universe leaves early time matter dominated decelerated phase for $\Gamma_{0}<1$ and ends with completely DE dominated accelerated phase (see lower left plot of Fig. 4).
    3. (c)
      $-1<w<0:$ In this interval of $w$, the system (31) provides four isolated critical points $B_{0},~{}B_{1},~{}B_{2}$ and $B_{3}$. In this situation, at late times, $z$ goes to $1$. Thus, $B_{2}$ is a global attractor. Here, one can deduce that $\Gamma_{0}<3$ for which, the nature of eigenvalues indicate that $B_{0}$ is saddle point, $B_{1}$ is unstable in $\Gamma_{0}<3$, and $B_{3}$ is saddle in $\Gamma_{0}<3$. The point $B_{0},~{}B_{2}$ correspond to accelerating epoch for $w<-1/3$ and $B_{1},~{}B_{3}$ represents accelerating epoch in $\Gamma_{0}>1$. If $w$ mimics the quintessential DE equation of state (i.e. $-1<w<-1/3$) with $\Gamma_{0}<1$, we obtain the proper evolution of the universe.
  2. 2.
    When $03w+\Gamma_{0}=$ 0: Now $x^{\prime}=0$, meaning that the trajectories are the straight lines $x=$ constant. The equation for $z$ becomes $z^{\prime}=\frac{3}{2}z(1-z)(1+w)$ and the decelerating parameter takes the form $q=(1+3w)/2$ which implies that all points in the phase space are accelerated for $w<-1/3$ and decelerated for $-1/3<w<0$.
    1. (a)
      $w<-1:$ In this phantom region of $w$, we have extracted two critical lines, namely, $B_{6}$, $B_{7}$, where DE and DM coexist. Here, $z$ is decreasing, deducing that for $w<-1$, $z$ goes to 00. Thus $B_{6}$ gives stable character and $B_{7}$ shows unstable qualitative behavior, which are verified using the eigenvalues (see Table 2).
    2. (b)
      $w=-1:$ When $w=-1$ i.e. $\Gamma_{0}=3$, we we have $z^{\prime}=0$, that is, all the points of the phase space are critical points and they are accelerated.
    3. (c)
@@
      $-1<w<0:$ Finally when $w>-1$, once again we have found two critical lines $\left(B_{6},B_{7}\right)$, having both DE and DM. From the evolution of $z$, one can conclude, $z$ is increasing i.e. $z$ goes to $1$. Therefore, $B_{7},~{}B_{6}$ behave like an attractor and a repeller respectively. These are also checked by eigenvalues (see Table 2). The phase plot is shown in the upper right plot of the Fig. 4.
  3. 3.
    When $03w+\Gamma_{0}>$ 0: In that case, at late times, $x$ goes to $1$, and, at late times, the equation for $z$ will become: $z^{\prime}=\frac{3}{2}z(1-z)\left(1-\frac{\Gamma_{0}}{3}\right)$. Here, it will appear the following situations:
    1. (a)
@@
      $0<\Gamma_{0}<3:$ Here, one can verify that $w>-1$. In that case, the dynamical system (31) produces four critical points $B_{0},~{}B_{1},~{}B_{2},~{}B_{3}$. As $z^{\prime}>0$, $z$ goes to $1$, and thus, $B_{3}$ is a global attractor. Again, $B_{0}$ is unstable and $B_{1},~{}B_{2}$ are saddle by behavior. If $-1/3<w$ and $\Gamma_{0}>1$, the evolution scenario is at early phase, orbits leaves from second fluid dominated decelerating phase and ends in DM dominated accelerating phase (see upper left plot of Fig. 4).
    2. (b)
      $\Gamma_{0}=3:$ Here, one can easily find two isolated critical points $B_{0},~{}B_{2}$ and one critical line $B_{5}$. Now, we will have $z^{\prime}=\frac{3}{2}z(1-z)(1+w)(1-x)$. Taking into account that, $w>-1$, we conclude that $B_{5}$ behaves like a global attractor. From the sign of eigenvalues, we can determine the stability of $B_{0},~{}B_{2}$ ($B_{0}$ is unstable and $B_{2}$ is saddle). Therefore, the final state of the universe is DM dominated accelerating phase.
    3. (c)
      $\Gamma_{0}>3:$ In that case, once again, there are four critical points $B_{0},~{}B_{1},~{}B_{2},~{}B_{3}$ in the phase space. Since, $z$ goes to 00, and thus, $B_{1}$ is a global attractor. From the point of view of the nature of eigenvalues, $B_{3}$ is always saddle, $B_{0}$ is unstable where $B_{2}$ is saddle, and $B_{2}$ is unstable where $B_{0}$ is saddle. Here, we can observe that $B_{0},~{}B_{2}$ for $w<-1/3$ and $B_{1},~{}B_{3}$ always, lie in the accelerating phase.
One fluid system
@@
Critical point Existence Stability $\Omega_{\rm dm}$ $q$ Acceleration
$B^{\prime}_{0}(1,0)$ Always Stable if $\Gamma_{0}>3$ and 1 $\frac{1}{2}(1-\Gamma_{0})$ $\Gamma_{0}>1$
@@
$B^{\prime}_{1}\left(1,1\right)$ Always Stable if $\Gamma_{0}<3$ and 1 $\frac{1}{2}(1-\Gamma_{0})$ $\Gamma_{0}>1$
@@
$B^{\prime}_{2}\left(1,z_{c}\right)$ $\Gamma_{0}=3$ All are critical points $1$ $-1$ Yes
Two fluid system
Critical point Existence Eigenvalue Stability $\Omega_{f}$ $\Omega_{\rm dm}$ $q$ Acceleration
$B_{0}(0,0)$ Always $\left(3w+\Gamma_{0},\frac{3}{2}(1+w)\right)$ Stable if $13w+\Gamma_{0}<0,~{}w\leq-$ 1 $1$ 00 $\frac{1}{2}(1+3w)$ $w<-\frac{1}{3}$
and if $13w+\Gamma_{0}=0,~{}w<-$ 1;
Otherwise Unstable or Saddle
$B_{1}(1,0)$ Always $\left(-(3w+\Gamma_{0}),\frac{3-\Gamma_{0}}{2}\right)$ Stable if $3w+\Gamma_{0}>0,~{}\Gamma_{0}\geq 3$ 00 $1$ $\frac{1}{2}(1-\Gamma_{0})$ $\Gamma_{0}>1$
and if $33w+\Gamma_{0}=0,~{}\Gamma_{0}>$ 3;
Otherwise Unstable or Saddle
$B_{2}(0,1)$ Always $\left(3w+\Gamma_{0},-\frac{3}{2}(1+w)\right)$ Stable if $13w+\Gamma_{0}<0,~{}w\geq-$ 1 $1$ 00 $\frac{1}{2}(1+3w)$ $w<-\frac{1}{3}$
and if $13w+\Gamma_{0}=0,~{}w>-$ 1;
Otherwise Unstable or Saddle
$B_{3}(1,1)$ Always $\left(-(3w+\Gamma_{0}),\frac{\Gamma_{0}-3}{2}\right)$ Stable if $3w+\Gamma_{0}>0,~{}\Gamma_{0}\leq 3$ 00 $1$ $\frac{1}{2}(1-\Gamma_{0})$ $\Gamma_{0}>1$
and if $33w+\Gamma_{0}=0,~{}\Gamma_{0}<$ 3;
Otherwise Unstable or Saddle
$B_{4}\left(0,z_{c}\right)$ $w=-1$ $\left(\Gamma_{0}-3,0\right)$ Stable if $\Gamma_{0}<3$ and $1$ 00 $-1$ Yes
Unstable if $\Gamma_{0}>3$
$B_{5}\left(1,z_{c}\right)$ $\Gamma_{0}=3$ $\left(-3(1+w),0\right)$ Stable if $w>-1$ and 00 $1$ $-1$ Yes
Unstable if $w<-1$
$B_{6}\left(x_{c},0\right)$ $03w+\Gamma_{0}=$ 0 $\left(0,\frac{3-\Gamma_{0}}{2}\right)$ Stable if $\Gamma_{0}>3$ and $1-x_{c}$ $x_{c}$ $\frac{1}{2}(1+3w)$ $w<-\frac{1}{3}$
Unstable if $\Gamma_{0}<3$
$B_{7}\left(x_{c},1\right)$ $03w+\Gamma_{0}=$ 0 $\left(0,\frac{\Gamma_{0}-3}{2}\right)$ Stable if $\Gamma_{0}<3$ and $1-x_{c}$ $x_{c}$ $\frac{1}{2}(1+3w)$ $w<-\frac{1}{3}$
Unstable if $\Gamma_{0}>3$
$B_{8}\left(x_{c},z_{c}\right)$ $w=-1,~{}\Gamma_{0}=3$ $(0,0)$ All points in the phase $1-x_{c}$ $x_{c}$ $-1$ Yes
@@
Table 2: Summary of the critical points, their existence, stability and the values of the cosmological parameters at those points for the dynamical system (30) and (31) with the matter creation rate $\Gamma=\Gamma_{0}H$.
Figure 3: Left Plot: This is the vector plot of the one dimensional system (30), i.e. for the model $\Gamma=\Gamma_{0}H$, considering positive $\Gamma_{0}$ with $\Gamma_{0}<3$. Right Plot: This is the vector plot of the one dimensional system (30) considering positive $\Gamma_{0}$ with $\Gamma_{0}>3$.
Figure 4: Description of the phase space of the two-fluid system for the matter creation rate $\Gamma=\Gamma_{0}H$. Upper Left Plot: The phase plot of the system (31) assuming $w=0.2$ and $\Gamma_{0}=2$. If we take any value of $w$ from $w>-1$ and $\Gamma_{0}$ in the region $30<\Gamma_{0}<$ 3, satisfying $03w+\Gamma_{0}>$ 0, we can get similar plot. Upper Right Plot: The phase plot of the system (31) for $-1<w<0$ satisfying $03w+\Gamma_{0}=$ 0. Here when we have assumed $w=-0.8$ and $\Gamma_{0}=2.4$, however, any value of $w$ lying in $(-1,0)$ and any positive value of $\Gamma_{0}$ satisfying the above condition will give similar plot. Lower Left Plot: The phase plot of the system (31) when the equation of state of the second fluid $w$ adopts the value $w=-1$. In this context, we choose $\Gamma_{0}=0.2$ but any positive value of $\Gamma_{0}$ satisfying $03w+\Gamma_{0}<$ 0 produces similar type of phase space diagram. Lower Right Plot: This is the phase plot of the system (31) for $w<-1$ satisfying $03w+\Gamma_{0}<$ 0. Here we use $w=-1.8$ and $\Gamma_{0}=0.1$ for drawing the graphics, however, any typical value of $w<-1$ and any positive value of $\Gamma_{0}$ $(\Gamma_{0}\neq 3)$ will give similar plot. Here the green region corresponds to the decelerating phase ($q>0$), cyan region represents the accelerating phase with $-1<q<0$ and the yellow region corresponds to the super accelerating phase ($q<-1$). In the lower right graph, the vertical black dotted line corresponds to $q=-1$.

### v2

@@ III.1 Model: $\Gamma=\Gamma_{0}$
When w<1w<-1, we again obtain five critical points which are shown in Table  1, i.e. $A_{0},~A_{1},~A_{2},~A_{3}$ and $A_{4}$. Since the eigenvalues are both negative at $A_{0}$, this is a stable point that exhibits dark-energy dominated accelerated universe. Being two eigenvalues, positive at $A_{1}$, this point is unstable which leads to the matter dominated decelerating phase of the universe. Clearly, the points $A_{0},~A_{2}$ properly describe the phase of the universe. Again, as eigenvalues are opposite in sign evaluated at $A_{3},~A_{4}$, these two critical points are saddle by nature representing acceleration with completely matter domination. Clearly, $A_{2}$, dark-energy dominated point is unstable because one can easily check that $x^{\prime}$, $z^{\prime}$ are both positive on $z=1$ and $x=0$, respectively. The lower right plot of Fig. 1 exhibits the phase space behavior where a class of orbits leave the decelerating phase and end in the super accelerating phase which is expected in the phantom cosmology, consistent with the trend shown in the lower right plot in Fig. 2. Here we notice that $w_{\rm dm}^{\rm eff}$ has a Gaussian like feature. In particularly, $w_{\rm dm}^{\rm eff}$ attains a minimum around the present epoch and then approaches 0 with the evolution of the universe.
[…]
@@ III.2 Model: $\Gamma=\Gamma_{0}H$
Here, we consider the particle creation rate, $\Gamma=\Gamma_{0}H$ where $\Gamma_{0}$ is a positive constant. In the following we analyze its effects on the universe evolution for the two-fluid system.
@@
where $\Gamma_{0}$ is a positive constant. The dynamical system is free from singularity. One can clearly see that the above system has four invariant manifolds x=0,x=1,z=0x=0,~x=1,~z=0 and $z=1$ which make the phase space domain $\bf R$}, a positively invariant set. In this model the decelerating parameter takes the form $q=\frac{1}{2}\left[1+3w(1-x)-\Gamma_{0}x\right]$ which does not contain the $z$ variable, and for $\left(1-\Gamma_{0}\right)x+(1+3w)(1-x)<0$, the accelerating expansion of the universe is realized. In the Table 2, we properly present the inherent property of the dynamical system (24) (i.e. critical points and different values of the cosmological parameters at these critical points). Since the evolution of $x$ variable significantly depends on the factor $\left(3w+\Gamma_{0}\right)$ for which we carry out the phase space analysis in three possible ways, such as:
@@
When $3w+\Gamma_{0}<0$, in this case, one can notice from Table 2 that there are four isolated critical points $B_{0},~B_{1},~B_{2},~B_{3}$ and two critical lines $B_{4}$(if w=1w=-1), $B_{5}$  (if $\Gamma_{0}=3$). Here, at late times, $x$ goes to zero, and then, at late times, we will have $z^{\prime}=\frac{3}{2}z(1-z)(1+w)$. Thus, the following sub-cases will appear: (a) when w<1,w<-1, in this case, we can see that, for $\Gamma_{0}\neq 3,$ the system has only four isolated critical points, namely, $B_{0},~B_{1},~B_{2},~B_{3}$. At late times, $z$ goes to zero, concluding that $B_{0}$ is a global attractor. Again, examining the eigenvalues, one can conclude that $B_{1}$ is saddle for $\Gamma_{0}>3$ and unstable for $\Gamma_{0}<3$, $B_{2}$ is always saddle in nature, $B_{3}$ is unstable if $\Gamma_{0}>3$ and saddle by behavior if $\Gamma_{0}<3$. Here, $B_{0},~B_{2}$ always lie in the accelerated region while $B_{1},~B_{3}$ remain in the accelerated phase if $\Gamma_{0}>1$. So we can conclude that $B_{0}$ is the late time DE dominated accelerating global attractor and universe undergoes from decelerating to accelerating phase if we choose the model parameter $\Gamma_{0}<1$. This scenario is depicted in the lower right plot of the Fig. 4 and in the right plot of the Fig. 5. On the other hand, for $\Gamma_{0}=3,$ along with two isolated critical points $B_{0},~B_{2}$, we have one critical line $B_{5}$, containing the critical points $B_{1},~B_{3}$. As one eigenvalue is positive and other is zero, $B_{5}$ behaves like unstable point. All the critical points and line are representing accelerating solutions. Once again, $B_{0}$ is the only late time DE dominated global attractor which we have seen in the previous paragraph. When the EoS of the second fluid coincides with cosmological constant (w=1)(w=-1), one can get two critical points and one critical line, namely, $B_{1},~B_{3}$ and $B_{4}$. Now, we have $z^{\prime}=\frac{3}{2}z(1-z)x\left(1-\frac{\Gamma_{0}}{3}\right)$. Taking into account that $1-\frac{\Gamma_{0}}{3}>0$, we deduce that $z$ is an increasing function, concluding that $B_{4}$ behaves like stable in nature. The line $B_{4}$ is always accelerated while $B_{1},~B_{3}$ are accelerated for $\Gamma_{0}>1$ otherwise decelerated. Thus, one may conclude that the universe leaves early time matter dominated decelerated phase for $\Gamma_{0}<1$ and ends with completely DE dominated accelerated phase (see lower left plot of Fig. 4 and middle plot of Fig. 5). When 1<w<0-1<w<0, in this interval of $w$, the system (24) provides four isolated critical points $B_{0},~B_{1},~B_{2}$ and $B_{3}$. In this situation, at late times, $z$ goes to $1$. Thus, $B_{2}$ is a global attractor. Here, one can deduce that $\Gamma_{0}<3$ for which, the nature of eigenvalues indicate that $B_{0}$ is saddle point, $B_{1}$ is unstable in $\Gamma_{0}<3$, and $B_{3}$ is saddle in $\Gamma_{0}<3$. The point $B_{0},~B_{2}$ correspond to accelerating epoch for w<1/3w<-1/3 and $B_{1},~B_{3}$ represents accelerating epoch in $\Gamma_{0}>1$. If $w$ mimics the quintessential DE EoS (i.e. 1<w<1/3-1<w<-1/3) with $\Gamma_{0}<1$, we obtain the proper evolution of the universe.
@@
When $3w+\Gamma_{0}=0$, $x^{\prime}=0$, meaning that the trajectories are the straight lines $x=$ constant. The equation for $z$ becomes $z^{\prime}=\frac{3}{2}z(1-z)(1+w)$ and the decelerating parameter takes the form $q=(1+3w)/2$ which implies that all points in the phase space are accelerated for w<1/3w<-1/3 and decelerated for 1/3<w<0-1/3<w<0. When w<1,w<-1, in this phantom region of $w$, we have extracted two critical lines, namely, $B_{6}$, $B_{7}$, where DE and DM coexist. Here, $z$ is decreasing, deducing that for w<1w<-1, $z$ goes to 0. Thus $B_{6}$ gives stable character and $B_{7}$ shows unstable qualitative behavior, which are verified using the eigenvalues (see Table 2). When w=1w=-1, i.e. $\Gamma_{0}=3$, we we have $z^{\prime}=0$, that is, all the points of the phase space are critical points and they are accelerated. When 1<w<0,-1<w<0, once again we have found two critical lines $\left(B_{6},B_{7}\right)$, having both DE and DM. From the evolution of $z$, one can conclude, $z$ is increasing i.e. $z$ goes to $1$. Therefore, $B_{7},~B_{6}$ behave like an attractor and a repeller respectively. These are also checked by eigenvalues (see Table 2). The phase plot is shown in the upper right plot of Fig. 4.
@@
When $3w+\Gamma_{0}>0$, in that case, at late times, $x$ goes to $1$, and, at late times, the equation for $z$ will become: $z^{\prime}=\frac{3}{2}z(1-z)\left(1-\frac{\Gamma_{0}}{3}\right)$. Here, it will appear the following situations, such as, $0<\Gamma_{0}<3:$ Here, one can verify that w>1w>-1. In that case, the dynamical system (24) produces four critical points $B_{0},~B_{1},~B_{2},~B_{3}$. As $^{\prime}>0, zz$ goes to $1$, and thus, $B_{3}$ is a global attractor. Again, $B_{0}$ is unstable and $B_{1},~B_{2}$ are saddle by behavior. If 1/3<w-1/3<w and $\Gamma_{0}>1$, the evolution scenario is at early phase, orbits leaves from second fluid dominated decelerating phase, reflecting the absence of a past DM dominated era, and ends in DM dominated accelerating phase (see upper left plot of Fig. 4 and left plot of Fig. 5). For $\Gamma_{0}=3$, one can easily find two isolated critical points $B_{0},~B_{2}$ and one critical line $B_{5}$. Now, we will have $z^{\prime}=\frac{3}{2}z(1-z)(1+w)(1-x)$. Taking into account that, w>1w>-1, we conclude that, $B_{5}$ behaves like a global attractor. From the sign of eigenvalues, we can determine the stability of $B_{0},~B_{2}$ ($B_{0}$ is unstable and $B_{2}$ is saddle). Therefore, the final state of the universe is DM dominated accelerating phase. For $\Gamma_{0}>3$, In that case, once again, there are four critical points $B_{0},~B_{1},~B_{2},~B_{3}$ in the phase space. Since, $z$ goes to 0, and thus, $B_{1}$ is a global attractor. From the point of view of the nature of eigenvalues, $B_{3}$ is always saddle, $B_{0}$ is unstable where $B_{2}$ is saddle, and $B_{2}$ is unstable where $B_{0}$ is saddle. Here, we can observe that $B_{0},~B_{2}$ for w<1/3w<-1/3 and $B_{1},~B_{3}$ always, lie in the accelerating phase.
For this model, the effective EoS of DM is given by $w_{\rm dm}^{\rm eff}=-\Gamma_{0}/3$, which remains a constant negative value throughout the evolution (see Fig. 5). For $\Gamma_{0}=2$, it behaves like dark energy, whereas for $\Gamma_{0}=0.2$ and $0.1$, it remains slightly negative but close to zero.
Fig. 6 shows the redshift evolution of the normalized Hubble parameter $H/H_{0}$ for the matter creation model $\Gamma=\Gamma_{0}H$, together with the standard $\Lambda{\rm CDM}$ prediction (upper panel). The fractional deviation, defined as $\Delta H/H=(H_{\rm Model}-H_{\Lambda{\rm CDM}})/H_{\Lambda{\rm CDM}}$, is plotted in the lower panel. We find that both models exhibit nearly identical behavior at low redshifts. At higher redshifts, however, the matter creation scenario departs significantly from $\Lambda{\rm CDM}$. These deviations illustrate how matter creation models can produce observationally testable signatures in the cosmic expansion history.
@@
Critical point Existence Eigenvalue Stability $\mathbf{\Omega_{f}}$ $\mathbf{\Omega_{\rm dm}^{\rm eff}}$ $\mathbf{q}$ Acceleration
$B_{0}(0,0)$ Always $\left(3w+\Gamma_{0},\frac{3}{2}(1+w)\right)$ Stable if $3w+\Gamma_{0}<0,~w\leq-1$ 11 0 $\frac{1}{2}(1+3w)$ $w<-\frac{1}{3}$
and if $3w+\Gamma_{0}=0,~w<-1$;
Otherwise Unstable or Saddle
$B_{1}(1,0)$ Always $\left(-(3w+\Gamma_{0}),\frac{3-\Gamma_{0}}{2}\right)$ Stable if $3w+\Gamma_{0}>0,~\Gamma_{0}\geq 3$ 0 11 $\frac{1}{2}(1-\Gamma_{0})$ $\Gamma_{0}>1$
and if $3w+\Gamma_{0}=0,~\Gamma_{0}>3$;
Otherwise Unstable or Saddle
$B_{2}(0,1)$ Always $\left(3w+\Gamma_{0},-\frac{3}{2}(1+w)\right)$ Stable if $3w+\Gamma_{0}<0,~w\geq-1$ 11 0 $\frac{1}{2}(1+3w)$ $w<-\frac{1}{3}$
and if $3w+\Gamma_{0}=0,~w>-1$;
Otherwise Unstable or Saddle
$B_{3}(1,1)$ Always $\left(-(3w+\Gamma_{0}),\frac{\Gamma_{0}-3}{2}\right)$ Stable if $3w+\Gamma_{0}>0,~\Gamma_{0}\leq 3$ 0 11 $\frac{1}{2}(1-\Gamma_{0})$ $\Gamma_{0}>1$
and if $3w+\Gamma_{0}=0,~\Gamma_{0}<3$;
Otherwise Unstable or Saddle
$B_{4}\left(0,z_{c}\right)$ w=1w=-1 $\left(\Gamma_{0}-3,0\right)$ Stable if $\Gamma_{0}<3$ and 11 0 1-1 Yes
Unstable if $\Gamma_{0}>3$
$B_{5}\left(1,z_{c}\right)$ $\Gamma_{0}=3$ $\left(-3(1+w),0\right)$ Stable if w>1w>-1 and 0 11 1-1 Yes
Unstable if w<1w<-1
$B_{6}\left(x_{c},0\right)$ $3w+\Gamma_{0}=0$ $\left(0,\frac{3-\Gamma_{0}}{2}\right)$ Stable if $\Gamma_{0}>3$ and $1-x_{c}$ $x_{c}$ $\frac{1}{2}(1+3w)$ $w<-\frac{1}{3}$
@@
$B_{7}\left(x_{c},1\right)$ $3w+\Gamma_{0}=0$ $\left(0,\frac{\Gamma_{0}-3}{2}\right)$ Stable if $\Gamma_{0}<3$ and $1-x_{c}$ $x_{c}$ $\frac{1}{2}(1+3w)$ $w<-\frac{1}{3}$
@@
$B_{8}\left(x_{c},z_{c}\right)$ $w=-1,~\Gamma_{0}=3$ $(0,0)$ All points in the phase $1-x_{c}$ $x_{c}$ 1-1 Yes
@@
Table 2: Summary of the critical points, their existence, stability and the values of the cosmological parameters at those points for the dynamical system (24) with the matter creation rate $\Gamma=\Gamma_{0}H$.
Figure 4: Description of the phase space of the two-fluid system for the matter creation rate $\Gamma=\Gamma_{0}H$. Upper Left Plot: The phase plot of the system (24) assuming $w=0.1$ and $\Gamma_{0}=2$. If we take any value of $w$ from w>1w>-1 and $\Gamma_{0}$ in the region $0<\Gamma_{0}<3$, satisfying $3w+\Gamma_{0}>0$, we can get similar plot. Upper Right Plot: The phase plot of the system (24) for 1<w<0-1<w<0 satisfying $3w+\Gamma_{0}=0$. Here when we have assumed w=0.8w=-0.8 and $\Gamma_{0}=2.4$, however, any value of $w$ lying in (1,0)(-1,0) and any positive value of $\Gamma_{0}$ satisfying the above condition will give similar plot. Lower Left Plot: The phase plot of the system (24) when the EoS of the second fluid $w$ adopts the value w=1w=-1. In this context, we choose $\Gamma_{0}=0.2$ but any positive value of $\Gamma_{0}$ satisfying $3w+\Gamma_{0}<0$ produces similar type of phase space diagram. Lower Right Plot: This is the phase plot of the system (24) for w<1w<-1 satisfying $3w+\Gamma_{0}<0$. Here we use w=1.8w=-1.8 and $\Gamma_{0}=0.1$ for drawing the graphics, however, any typical value of w<1w<-1 and any positive value of $\Gamma_{0}$ $(\Gamma_{0}\neq 3)$ will give similar plot. Here the green region corresponds to the decelerating phase ($q>0$), cyan region represents the accelerating phase with 1<q<0-1<q<0 and the yellow region corresponds to the super accelerating phase (q<1q<-1). In the lower right graph, the vertical black dotted line corresponds to q=1q=-1.
Figure 5: We display the evolution plots of the DM density parameter $\Omega_{\rm dm}^{\rm eff}$, the second fluid density parameter $\Omega_{f}$, the deceleration parameter $q$, and the effective equation of state of DM $w_{\rm dm}^{\rm eff}$ corresponding to the dynamical system (24), with the matter creation rate $\Gamma=\Gamma_{0}H$. Left Plot: We have drawn this plot using the model parameters $w=0.1$ and $\Gamma_{0}=2$. Here, the trajectories end in DM dominated accelerated quintessence era. In this case, the past DM-dominated era can not be realized since no decelerating DM-dominated saddle point appears (see the upper left plot of Fig. 4). Middle Plot: This plot is produced adopting the values of the model parameters w=1w=-1 and $\Gamma_{0}=0.2$ where the late time evolution of the universe is attracted by DE dominated cosmological constant era connecting through DM dominated decelerated era. Right Plot: For this plot, we take w=1.2w=-1.2 and $\Gamma_{0}=0.1$ where one can see that trajectories are attracted by DE dominated phantom era at late-times joining through a DM dominated decelerated phase.
[…]